OpenStudy (anonymous):

Find the dementions (radius and height) of a cylindrical can in order to minimize the surface area and have a fixed volume of 1000 cubic centimeters help please i dont get it

OpenStudy (anonymous):

do you have any idea how to do this?

OpenStudy (ybarrap):

yes, do you know the Method of La Grange?

OpenStudy (anonymous):

No i dont, my math teacher never mentioned that.

OpenStudy (ybarrap):

Ok. l'll show you by using it.

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

?

OpenStudy (ybarrap):

A=surface area = 2*pi*r^2 + 2pi*r*h, including ends of can and sides Constraint: V = volume = pi*r^2*h = 1000 cm^2 dA/r = 4*pi*r + 2*pi*h= L * dV/r = L * 2pi*r*h (eq 1) dA/h = 2*pi*r = L * 2*r^2 (eq 2) Need to first find h and r in terms of L then plug into constraint to find minimized surface area. from eq 1: 2*pi*r = L * 2 * r^2 2*pi/(2L) = r from eq 2: 4*pi*r + 2*pi*h = L * 2pi*r*h 4*pi*r = L * 2pi*r*h - 2pi*h = h(L*2*pi*r - 2*pi) h = 4*pi*r/(L*2pi*r - 2*pi) = 4*pi*r/( 2*pi (L*r - 1) ) = 2*r/(L*r - 1) = 2*(2*pi/(2L))/(L*r - 1) = 8*pi*L/(L*(2*pi/(2L)) - 1) = 8*pi*L/(pi - 1) Using constraint: pi*r^2*h = 1000 pi*(pi/L)^2*8*pi*L/(pi - 1) = 1000 8Lpi^3/(pi - 1) = 1000 L = (pi - 1)*1000/(8*pi^3) = (pi - 1)*250/(4*pi^3) = 4.32 (approx) Plug back into r and h: 2*pi/(2L) = r r = pi/L = pi/4.32 = 0.73 h = 8*pi*L/(pi - 1) = 50.70 Plug into surface area (A) to find minimized value: A = 2*pi*r^2 + 2pi*r*h = 2*pi*0.73^2 + 2pi*0.73*50.70 = 236.00 Ans. 236.00 cm^2 Please check my arithmetic, but the essence of The La Grange is all here. Easy process but lots of calculations.

OpenStudy (anonymous):

but isnt it suppose to be cm^3 not cm ^2?

OpenStudy (ybarrap):

Area is square cm and volume is cubic, because area is 2 dim and volume is 3 dim. Also, question asks for r and h not area. I also have an error in the first part, I'll correct below.

OpenStudy (anonymous):

so can you list what the radius and height will be? i dont know where the 236.00 belongs hehe

OpenStudy (ybarrap):

A=surface area = 2*pi*r^2 + 2pi*r*h, including ends of can and sides Constraint: V = volume = pi*r^2*h = 1000 cm^2 dA/r = 4*pi*r + 2*pi*h= L * dV/r = L * 2pi*r*h (eq 1) dA/h = 2*pi*r = L * dV/h = L * pi*r^2 (eq 2) Need to first find h and r in terms of L then plug into constraint to find minimized surface area. from eq 2: 2*pi*r = L * pi * r^2 2/L = r from eq 1: 4*pi*r + 2*pi*h = L * 2pi*r*h 4*pi*r = L * 2pi*r*h - 2pi*h = h(L*2*pi*r - 2*pi) h = 4*pi*r/(L*2pi*r - 2*pi) = 4*pi*r/( 2*pi (L*r - 1) ) = 2*r/(L*r - 1) = 2*(2/L))/(L*r - 1) = 4/(L*(2/L)) - 1) = 4/(2 - 1/L) Using constraint: pi*r^2*h = 1000 pi*(2/L)^2*4/(2 - 1/L) = 1000 16*pi/(2L - 1) = 1000 16*pi/1000 = 2L - 1 2L = 4pi/250 + 1 L = 2pi/250 + 1 L = 1.03 (approx) Plug back into r and h: 2/L = r r = 2/1.03 = 1.95 cm h = 4/(2 - 1/L) = 3.87 cm Plug into surface area (A) to find minimized value: A = 2*pi*r^2 + 2pi*r*h = 2*pi*1.95^2 + 2pi*1.95*3.87 = 71.31 cm^2 Ans. r = 1.95 cm, h = 3.87 cm Please check my arithmetic, but the essence of The La Grange is all here. Easy process but lots of calculations.

OpenStudy (ybarrap):

To see if area is minimum, try another value of r and solve for h in constraint. Area should be larger.

OpenStudy (anonymous):

is there any way you can help me with the other question i asked? i promise its not this much work hehe.